Circular No. 7973

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002fj IN NGC 2642 L. A. G. Monard, Bronberg Ridge (southeast of Pretoria), S. Africa,

reports his discovery of an apparent supernova (mag 15.8) on three unfiltered CCD images taken with a 0.30-m reflector around Sept. 12.12 UT; additional unfiltered magnitudes from Monard: Sept. 13.12, 15.6; 14.122, 15.5; 15.108, 15.5; 16.122, 15.6; 17.118, 15.9; 18.114, 16.2. SN 2002fj is located at α = 8^{h}40^{m}45^{s}.10, δ = −4^{o}07′38′′.5 (equinox 2002.0), which is 12′′ east and 21′′.5 south of the center of NGC 2642 and situated near the end of the bar and at the beginning of a spiral arm. Nothing appears at this location on an image taken by Monard in June (limiting mag 17) or on Digitized Sky Survey images. A bright foreground star is located just to the north of the new object.

SUPERNOVA 2002fk IN NGC 1309 Reports of the independent discoveries on unfiltered CCD images of

an apparent supernova in NGC 1309 have been received from R. Kushida, Yatsugatake South Base Observatory (via S. Nakano, Sumoto, Japan), and from J. Wang and Y. L. Qiu (on behalf of Beijing Astronomical Observatory supernova survey). A precise position for the new object has been measured by Y. Kushida from R. Kushida’s image of Sept. 17.719 UT: α = 3^{h}22^{m}05^{s}.71, δ = −15^{o}24′03′′.2 (equinox 2000.0), which is ∼ 12′′ west and 3′′.5 south of the center of NGC 1309 (offset somewhat uncertain due to difficulty in defining the nucleus center). Magnitude estimates (K = Kushida, W = Wang): 2000 Dec. 17, [18 (K); 2002 Feb. 6.50, [18.0: (W); Sept. 15.84, 15.7: (W); 17.719, 15.0 (K); 17.79, 14.5: (W).

V4742 SAGITTARII T. Kato, Kyoto University, reports the following prediscovery limiting

magnitudes by K. Haseda (Aichi, Japan, 0.10-m f/4 twin patrol camera) from T-Max 400 films: Feb. 23.490 UT, [12.8; June 18.627, [12.8. CCD photometry by J. D. West, Mulvane, KS: Sept. 17.095, B = 9.58; 17.101, R_{c} = 7.09; 17.117, V = 7.96; 18.079, V = 8.31, R_{c} = 7.37. Additional visual magnitude estimates (cf. IAUC 7972): Sept. 17.802, 8.5 (A. Kammerer, Ettlingen, Germany); 18.260, 8.4 (J. Bedient, Honolulu, HI); 18.508, 8.8 (A. Pearce, Nedlands, W. Australia).

2002 September 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green